# HarmoQ: Harmonized Post-Training Quantization for High-Fidelity Image Super-Resolution

### v1

@@ Introduction
Figure 3: Structural Residual Calibration analysis in SwinIR. (a) Optimal weight calibration $\Delta W^{*}$ heatmap using Laplacian filter, revealing structured channel relationships. (b) Response comparison before/after $\Delta W^{*}$ calibration showing structure-aware modulation. (c) Efficacy across layers for different projection matrices. Laplacian filter achieves the best performance (see Table 3 for quantitative comparison), while random projection fails due to lack of structure-aware design. Cost analysis via Frobenius norm of $\Delta W$ shows consistent magnitude across projection types.
@@
To validate our structural residual calibration, we analyze the optimal weight calibration $\Delta W^{*}$ across different projection matrices. Figure 3(a) shows the heatmap of $\Delta W^{*}$ derived via a Laplacian filter, revealing structured, non-random patterns that target specific channel relationships, confirming the structure-specific corrective capability of our method. Figure 3(b) illustrates how $\Delta W^{*}$ modulates activation responses while preserving key signal characteristics. The smooth transitions in the corrected profile support our structure-aware calibration principle and demonstrate reduced structural distortion. Figure 3(c) further shows consistent error reduction across layers for various structural projection filters. DCT high-pass and learned basis filters yield superior results, especially in deeper layers where quantization errors accumulate, while random projections perform poorly, underscoring the importance of a structure-aware design.
@@
Set5 ($\times 2$) Set14 ($\times 2$) BSD100 ($\times 2$) Urban100 ($\times 2$) Set5 ($\times 4$) Set14 ($\times 4$) BSD100 ($\times 4$) Urban100 ($\times 4$)
Method Bit PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
@@
Table 2: Component-wise ablation analysis.
Components Set5 ($\times 2$) Urban100 ($\times 4$)
SRC HSO ABR Config PSNR SSIM PSNR SSIM
@@
We evaluate HarmoQ on standard super-resolution benchmarks using SwinIR-light (Liang et al. 2021) and HAT-S (Chen et al. 2023a) as backbone networks. Following common practice, we train on DF2K dataset (combination of DIV2K (Timofte et al. 2017) and Flickr2K (Lim et al. 2017)) and test on five widely-used benchmarks: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde et al. 2010), B100 (Martin et al. 2001), Urban100 (Huang et al. 2015), and Manga109 (Matsui et al. 2017). We conduct experiments with ×2 and ×4 upscaling factors under 2-bit and 3-bit quantization settings. Performance is measured using PSNR and SSIM computed on the Y channel of YCbCr color space.
@@
Table 3: Structural feature projection matrix analysis.
Set5 ($\times 2$) Urban100 ($\times 4$)
Projection Matrix $H$ PSNR SSIM PSNR SSIM
@@
Our ablation study in Table 2 demonstrates that HarmoQ’s performance gains arise from complex nonlinear synergies rather than linear superposition. The baseline method, which processes weight and activation quantization in isolation while ignoring their coupling, embodies the dilemma of traditional super-resolution quantization. In contrast, our components show significant advantages. The SRC mechanism markedly improves performance by compensating for loss of structural information from activation quantization. HSO outperforms SRC when applied alone, highlighting the critical role of balancing quantization difficulty to enhance system robustness. ABR shows limited standalone effectiveness, as its performance depends on an established quantization balance. Pairwise combinations reveal strong synergy between SRC and HSO, where calibration is amplified by the quantization balance. Without HSO as a bridge, however, SRC and ABR struggle to synergize. Ultimately, the full HarmoQ framework outperforms all two-component combinations, demonstrating that its unified optimization enables a qualitative leap from local to global coordination. Consistent results on the Urban100 dataset validate this universality in scenes with rich textures.
@@
Our ablation study in Table 2 explores the quest for optimal structure-aware calibration. Laplacian filters emerge victorious (37.98/0.9602), their second-order nature perfectly capturing edge discontinuities that activation quantization disrupts. Sobel filters (37.85/0.9589) and learned basis (37.91/0.9594) engage in a close contest, proving both classical gradient wisdom and data-driven intelligence can decode quantization artifacts. DCT high-pass filters (37.72/0.9575) show promise yet fall short, suggesting spatial gradients trump pure frequency decomposition. Random projection’s failure (37.34/0.9548) delivers the final verdict: effective error compensation demands explicit structural semantics, not blind dimensionality reduction.

### v2

@@ Introduction
Figure 3: Structural Residual Calibration analysis in SwinIR. The top-left panel presents the optimal weight calibration $\Delta W^{*}$ heatmap using Laplacian filter, revealing structured channel relationships. The top-right panel shows response comparison before and after $\Delta W^{*}$ calibration, demonstrating structure-aware modulation. The bottom panel illustrates the efficacy across layers for different projection matrices. Laplacian filter achieves superior performance (see Table 3 for quantitative comparison), while random projection fails due to lack of structure-aware design. Cost analysis via Frobenius norm of $\Delta W$ shows consistent magnitude across projection types.
@@
To validate our structural residual calibration, we analyze the optimal weight calibration $\Delta W^{*}$ across different projection matrices. The top-left panel of Figure 3 shows the heatmap of $\Delta W^{*}$ derived via a Laplacian filter, revealing structured, non-random patterns that target specific channel relationships, confirming the structure-specific corrective capability of our method. The top-right panel illustrates how $\Delta W^{*}$ modulates activation responses while preserving key signal characteristics. The smooth transitions in the corrected profile support our structure-aware calibration principle and demonstrate reduced structural distortion. The bottom panel further shows consistent error reduction across layers for various structural projection filters. DCT high-pass and learned basis filters yield superior results, especially in deeper layers where quantization errors accumulate, while random projections perform poorly, underscoring the importance of a structure-aware design.
@@
[…]
@@
[…]
@@
We evaluate HarmoQ on standard super-resolution benchmarks using SwinIR-light (Liang et al. 2021) and HAT-S (Chen et al. 2023a) as backbone networks. Following common practice, we train on DF2K dataset (combination of DIV2K (Timofte et al. 2017) and Flickr2K (Lim et al. 2017)) and test on five widely-used benchmarks: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde, Elad, and Protter 2010), B100 (Martin et al. 2001), Urban100 (Huang, Singh, and Ahuja 2015), and Manga109 (Matsui et al. 2017). We conduct experiments with ×2 and ×4 upscaling factors under 2-bit and 3-bit quantization settings. Performance is measured using PSNR and SSIM computed on the Y channel of YCbCr color space.
@@
[…]
@@
Table 3: Structural feature projection matrix analysis
@@
Our ablation study in Table Baseline Methods. demonstrates that HarmoQ’s performance gains arise from complex nonlinear synergies rather than linear superposition. The baseline method, which processes weight and activation quantization in isolation while ignoring their coupling, embodies the dilemma of traditional super-resolution quantization. In contrast, our components show significant advantages. The SRC mechanism markedly improves performance by compensating for loss of structural information from activation quantization. HSO outperforms SRC when applied alone, highlighting the critical role of balancing quantization difficulty to enhance system robustness. ABR shows limited standalone effectiveness, as its performance depends on an established quantization balance. Pairwise combinations reveal strong synergy between SRC and HSO, where calibration is amplified by the quantization balance. Without HSO as a bridge, however, SRC and ABR struggle to synergize. Ultimately, the full HarmoQ framework outperforms all two-component combinations, demonstrating that its unified optimization enables a qualitative leap from local to global coordination. Consistent results on the Urban100 dataset validate this universality in scenes with rich textures.
@@
Our ablation study in Table Baseline Methods. explores the quest for optimal structure-aware calibration. Laplacian filters emerge victorious (37.98/0.9602), their second-order nature perfectly capturing edge discontinuities that activation quantization disrupts. Sobel filters (37.85/0.9589) and learned basis (37.91/0.9594) engage in a close contest, proving both classical gradient wisdom and data-driven intelligence can decode quantization artifacts. DCT high-pass filters (37.72/0.9575) show promise yet fall short, suggesting spatial gradients trump pure frequency decomposition. Random projection’s failure (37.34/0.9548) delivers the final verdict: effective error compensation demands explicit structural semantics, not blind dimensionality reduction.